You have 32 coins. They are all nickels and dimes. You have a total of 2.85. Write a system of equations to determine how many dimes and how many nickels you have
Answer by Maths68(1474)


Total Amount = $2.85=2.85*100 = 285 cents
Total Coins = 32
Total Dimes = x
One Dime = d =10 cents
Total Nickels = 32-x
One Nickel = n =5 cents
d*(x)+n*(32-x)=Total Amount
10*x+5*(32-x)=285
10x+160-5x=285
5x+160=285
5x=285-160
5x=125
5x/5=125/5
x=25


Total Dimes = x = 25
Total Nickels = 32-x = 32-25 = 7

Check
=====
=25*10 + 7*5
=250+35
=285